Subject: APO 3.5-m Users Committee minutes 9/19/05
From: Bruce Gillespie
Submitted: Tue, 18 Oct 2005 19:53:48 -0600
Message number: 929
(previous: 928,
next: 930
up: Index)
APO 3.5-m Users Committee Phonecon, 10/17/05
Attending: Suzanne Hawley, Michael Strauss, Karl Glazebrook, Bruce
Gillespie, John Bally, Fred Hearty, Rene Walterbos, John Wilson
(representing UVa as an observer), Jon Holtzman, Bruce Balick
Absent: Russet McMillan, Don York, Al Harper
Minutes taken by Bruce Gillespie
**********************************
Suzanne started the meeting by going over the action items from
previous meetings:
[previous months]
ACTION: All committee members should identify who their local
systems administrator (i.e. the person who would be in charge of
installing TUI and providing information on it to users) is so we can
contact them. Please send the name and contact information to Bruce
Gillespie (gillespi at apo.nmsu.edu) - STATUS: completed for
UWashington, Colorado, NMSU. Bruce will ask the committee members
from JHU, Princeton, Chicago, and UVa to provide theirs.
ACTION: Suzanne asked if NAIC will satisfy Nancy [Chanover]'s needs
for planetary observations, in lieu of certain special filters that
were not possible to fit into NIC-FPS. Fred said he thought that
some of the filters and short-exposure modes in NIC-FPS do fit the
bill for planetary work, but he agreed to talk to Nancy to be sure -
STATUS: not completed, Jon Holtzman will talk to Nancy.
[new actions from last month]
ACTION: Users Committee members need to inform their users that any
remaining Remark holdouts must notify Suzanne and explain why they
cannot migrate to TUI - STATUS: no comments were received, so Suzanne
will declare in an e-mail to apo35general that Remark will not be
available for use from here on out.
ACTION: BruceG agreed to send to the Users Committee the
constituents of these [Futures Committee] sub-groups - STATUS:
completed, see below.
1) Planet group: Phil Armitage (CU); Eric Agol (UW); Michael
Strauss (Princeton); Holland Ford (JHU); Andy Davis (UC); Kurt
Anderson (NMSU)
2) NIR spectroscopy: Mike Strutskie, John Wilson (UVa), and
Steve Majewski (UVa); Tom Harrison (NMSU); Al Harper (UC); Jill Knapp
(PU); John Bally (CU); Scott Anderson (UW); Rosie Wyse (JHU)
3) Multislit spectroscopy: Ed Turner (PU); Bob O'Connell (UVa);
Rene Walterbos, Jon Holtzman (NMSU); John Stocke (CU)
ACTION: BruceG will send the GRB URL to the Users Committee to share
with their users - STATUS: completed.
ACTION: Send these [2005 science highlights] to Suzanne and BruceG
not later than 1 November, and preferably sooner, so that your
institution will be fairly represented in the report. Also, send all
your 3.5-m-based publications and theses during the past two years to
Rene Walterbos, also not later than 1 November - STATUS: not
completed. Please submit reports by 1 November at the latest.
**********************************
User feedback, comments from Institutional representatives:
Princeton (Michael Strauss)--Michael had nothing much to report from
the users; everyone seems "happy".
Colorado (John Bally)--John said his recent class expedition to APO
was a "fabulous visit, although the weather was bad."
NMSU (Jon Holtzman)--Jon reported that there a lot of complaints
about the recent bad weather.
Johns Hopkins (Karl Glazebrook)--Karl had nothing to report, but
asked if the various institution's systems administrators could be
added to the APO website.
ACTION: Once we collect the the names of the designated
institutional systems administrators, BruceG will have them listed on
the APO web pages for reference.
UWashington (Bruce Balick)--BruceB had no issues to report.
**********************************
Discussion of telescope/instruments report:
BruceG summarized highlights from the report (see below). Fred
Hearty elaborated on NIC-FPS status, and highlighted the upgrades and
improvements accomplished during the summer shutdown, which included
fixes to the filter wheels, correlated and Fowler sampling, noise
reduction (although one source of noise still remains to be tracked
down and fixed), and the installation of a grism (warm slit) with
R~100s through the H-band. There is some order overlap in the
spectrum, so he recommends using the Y filter to suppress the
overlap, which gives relatively clean YJH spectral coverage. The
dewar vacuum problem was not solved by adding a new ion pump to the
system, so the adsorbers were put back in the dewar as a stopgap.
The fill fitting and bellows for the LN2 are suspected as the source
of the vacuum leak, and further work is underway to fix it. The
instrument is down at the moment, and is expected to be back in
service by the end of this week after new vacuum seals are installed.
[Below is the monthly telescope/instrument attachment submitted to
the committee prior to the meeting]:
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 9/16
through 10/14/05
Bruce Gillespie
0) Overview
We hosted a large group of undergraduate students from CU, led by
John Bally. Also, a group of UW donors visited site with Bruce
Balick as host. CorMASS is currently being reintegrated with the
telescope for its upcoming run. We had continuing problems with a
NIC-FPS vacuum leak, resulting in extended trouble-shooting site
visit by Fred Hearty. The problem is still unresolved, but the
instrument has/is been/being used for science for a few nights this
month. We finished on-sky engineering tuning of telescope after
summer shutdown work, and its performance is nominal. Other
instruments and systems are working fine. Weather was unusually wet
and cloudy for this time of year. The 4th Observing Specialist
position recruitment was started.
1) Telescope
Working nominally, only problem left is infrequent instrument rotator
over-current events, which is being addressed during engineering time
this coming week.
2) Instruments
o NIC-FPS: a serious vacuum leak in the instrument dewar
developed on the day after the first night of resumption of science
use of NIC-FPS, taking it out of service. Fred Hearty returned to
the site to diagnose and fix the problem, and after a couple of weeks
of tests, he was not able to find the source of the problem. But he
was able to modify the instrument vacuum system to the point it could
be used for science, which is the state it is in now. Further tests
and rework are under planned for the coming week. The instrument is
lightly scheduled over the coming weeks.
o DIS: The blue-medium grating suffered minor damage during
shipment to/from JHU, where it had been sent for throughput
measurements. An insurance claim was filed with the shipper for the
replacement cost, which has already been paid. The grating is seldom
used, so no serious operational impact is likely.
3) Engineering and CIF projects
A new version of TUI was released, which handles some of the new
features in NIC-FPS, and also resolved some general guiding issues.
A plan for completing the new drive-box project was developed, and
the detailed design of the new top-end project was resumed in
earnest, having been slowed by manpower being diverted to the summer
shutdown work.
4) Miscellaneous
4th Observing Specialist position recruitment started.
**********************************
UVa and CorMASS:
JohnW said that the UVa astronomers and students are excited about
the prospects of joining ARC, and are planning to have Triplespec at
the telescope in Oct. '06, and officially joining ARC in Jan. '07.
CorMASS is back at APO and has been re-commissioned for use this
quarter and into next year. There are two scheduled CorMASS runs in
November, and the instrument will stay in stand-by ready mode for the
foreseeable future. Suzanne mentioned that the next quarter
allotments will mention CorMASS availability. There was a question
how CorMASS spectra compare to NIC-FPS grism spectra, and Fred Hearty
said he'd look at this, but he mentioned that the principal purpose
of the NIC-FPS grism capability was to get quick NIR spectra of GRBs.
**********************************
Triplespec review:
Suzanne said that the Triplespec Final Design Review is scheduled to
take place on 6 - 7 December at APO. Alan Uomoto (Magellan) has
agreed to chair the review, with Tom O'Brien (OSU) and Al Harper
(UChicago) as reviewers, as well as Fred Hearty, Stephane Beland, and
others from the ARC engineering staff participating.
**********************************
GRB, APOLLO updates:
Suzanne reported that there have been no GRB alerts observed at APO
lately, largely because NIC-FPS has been down. For APOLLO, BruceG
said that there is an agreement with Tom Murphy to provide support to
APO that would partially enable the recruitment of the 4th Observing
Specialist.
**********************************
Future CIF projects:
For the Futures Committee, and for our planning purposes for the next
three years, Suzanne and Bruce have assembled a list of current and
future CIF projects, which is given below. Suzanne asked the Users
Committee members to pay particular attention to the list of
instrumentation proposals from the perspective of their constituents'
interests. The list is:
ARC 3.5m CIF Projects during 2005-2008
1. Projects already started and should be completed prior to June, 2008
A. Telescope
new top end (truss and 2ndary motions, safeties)
new drive boxes
TUI maintenance/upgrades
B. Instruments
NIC-FPS upgrades (not new detector)
Triplespec (design review, funding for detector)
DIS upgrades/fixes (JHU trade study, new red detector)
2. Projects not yet started, but under discussion. Some may
be completed prior to June, 2008.
A. Telescope
drive controllers and servo replacements
tertiary mount and actuators rebuild
autofocus
atmospheric dispersion correction
new direct drives
enclosure rotation servo upgrades
B. Instruments
SPIcam focal reducer
Goddard Fabry-Perot upgrades
UW high speed camera
NIC-FPS upgrades (new detector)
Instrument provided by UFlorida
**********************************
Need for back-up programs:
The group discussed the importance of having it be the responsibility
of each telescope user to have back-up program(s) when they observe
at the 3.5-m telescope. Preferably ones that are TAC-approved and
using a different instrument, these back-up programs can be executed
on short notice when an instrument malfunctions or the weather is
marginal. The committee members were asked by Suzanne to talk to
their users about this, and it was decided that reminding users of
the importance of this on a routine basis would be a good thing to
do. [ACTION: BruceG will write a paragraph about the importance of
having back-up programs, to be added to the e-mail automated message
we send observers prior to their runs]
**********************************
Board of Governors annual meeting preparations:
Rene said that this year's Board meeting should be pretty
straightforward, and that he is starting to receive publication and
thesis lists to include in the annual report. Suzanne also re-
emphasized the need for science highlights from the respective
institutions for inclusion in the report.
**********************************
Next Users Committee phonecon will be on Monday, 28 November, at 8:30
AM Pacific Time. Agenda and other materials will be sent to the
committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 929 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO